Shift report 20020103 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Jan 03 2002 - 00:37:27 EST

  • Next message: Apache: "Shift report 20020103 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Jan  3 00:37:27 2002, Shift: 16:00-24:00
    Supervisor: sjs, with: 
    Log book#11 pages 141 - 141
    
    Shift Summary:
    ------------------------------------------------------------
    Continuous store through most of run
    
    data runs 6249,6250,6251
    trigger 6 lost durring run 6250 and only discovered
    after run 6251 went for some time and no trigger 6
    found in trigger sum.  Problem traced to bad cable
    connection in feh--fixed by reseating cables.
    
    removed x10 attenuation from ptq's for si1_1, si1_2,
    si2_5, si3_1  (2 ptq boxes...)
    
    no new detector problems found.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Jan 03 2002 - 00:38:14 EST